# From Seeing to Doing: Bridging Reasoning and Decision for Robotic Manipulation

### v1

@@ Abstract
Achieving generalization in robotic manipulation remains a critical challenge, particularly for unseen scenarios and novel tasks. Current Vision-Language-Action (VLA) models, while building on top of general Vision-Language Models (VLMs), still fall short of achieving robust zero-shot performance due to the scarcity and heterogeneity prevalent in embodied datasets. To address these limitations, we propose FSD (From Seeing to Doing), a novel vision-language model that generates intermediate representations through spatial relationship reasoning, providing fine-grained guidance for robotic manipulation. Our approach combines a hierarchical data pipeline for training with a self-consistency mechanism that aligns spatial coordinates with visual signals. Through extensive experiments, we comprehensively validated FSD’s capabilities in both “seeing” and “doing”, achieving outstanding performance across 8 benchmarks for general spatial reasoning and embodied reference abilities, as well as on our proposed more challenging benchmark VABench. We also verified zero-shot capabilities in robot manipulation, demonstrating significant performance improvements over baseline methods in both SimplerEnv and real robot settings. Experimental results show that FSD achieves 54.1% success rate in SimplerEnv and 72% success rate across 8 real-world tasks, outperforming the strongest baseline by 30%. More visualizations are available on website.
@@ 1 Introduction
A driving force behind robotics research is the pursuit of generalization: creating agents capable of versatile action across diverse robotic platforms, extending beyond familiar tasks, objects, and environments while adapting to dynamic visual inputs. Current approaches [21, 3] leverage pre-trained Vision-Language Models (VLMs) and transform them into Vision-Language-Action Models (VLAs) using large-scale embodied datasets. This enables systems to interpret natural language instructions and generate robotic manipulation actions [2]. The intention is to capitalize on the generalization capabilities of VLMs pre-trained on internet-scale data, with the hope that resulting VLAs will adapt to novel scenarios involving unseen objects and tasks. However, empirical evidence [72, 67, 35] demonstrates that directly applying the generalization power of VLMs falls short of achieving strong zero-shot performance on completely novel tasks.
@@
We attribute the limited generalization in VLA-based systems to scarcity and heterogeneity of datasets. Robotics data remains limited compared to language and vision datasets, preventing similar scaling laws [18, 29]. Despite growth in embodied datasets [40], their insufficient coverage and diversity prevent robust zero-shot generalization. Additionally, robotic embodiment heterogeneity [54] causes significant variation in VLA outputs, making end-to-end supervised learning from vision and language to diverse action outputs a potentially unrealistic path toward generalization.
@@
We introduce a novel pipeline addressing generalization challenges in robotic manipulation. Our approach leverages VLMs’ visual understanding capabilities, augmented with step-by-step reasoning to extract unified mid-level structural representations independent of robot embodiment. This representation is key to generalizing learning across diverse physical interactions and dynamic behaviors. Specifically, this mid-level representation includes spatial affordance boxes/points and visual traces, each represented as marked coordinates within visual images. These visual aids provide expressive yet compact spatial information that enables effective reasoning and decisions, overcoming both scarcity and heterogeneity limitations. We introduce FSD (From Seeing to Doing), a model that generates visual intermediate representations through spatial understanding and reasoning (Fig. 1). The FSD model comprises three key components: (1) Spatial Relationship-Focused Visual Chain-of-Thought (SrCoT), which conducts multi-step reasoning anchored by object coordinates and spatial relationships, treating visual aid generation as a reasoning process; (2) A hierarchical data pipeline combining large-scale embodied datasets with common sense data, establishing a weak-to-strong capability enhancement training process; and (3) A self-consistency mechanism that aligns understanding and generation by binding spatial coordinates with specific visual signals. We employ supervised fine-tuning of VLMs using annotated spatial COT datasets for visual aids generation, then action executors follow spatial affordances and visual traces through simple planning methods to complete action execution.
@@
Our FSD model generalizes effectively to new instructions, and scenes through enhanced reasoning abilities. Our contributions include: (1) A novel paradigm bridging VLM reasoning with embodied decisions via visual aids; (2) The SrCoT method enabling multi-step reasoning for visual aid generation and guiding zero-shot manipulation; (3) Our weak-to-strong spatial reasoning and visual aids datasets, along with VABench, a manually annotated challenging benchmark for visual aids generation; (4) Superior performance across 8 benchmarks in spatial reasoning, free space reference, and visual aids generation, with zero-shot deployment achieving 54.1% success in SimplerEnv and 73% in 8 real-world tasks, outperforming mainstream VLA models by 30%.
@@ 2 Related Work
Spatial Understanding and Reasoning with VLMs Spatial understanding and reasoning [32, 48, 8, 28, 44] require VLMs to infer spatial information beyond 2D RGB images, a capability crucial for embodied AI applications such as navigation [48, 12, 22] and manipulation [66]. Recent advances include SpatialVLM [5], which explicitly incorporates spatial primitives and coordinate systems to enhance geometric reasoning capabilities. Similarly, SpatialRGPT [7] and SpatialBot [4] improve spatial capabilities through more precise spatial relationship modeling. FSD specifically targets embodied manipulation scenarios by enhancing spatial reasoning capabilities through novel SrCoT mechanisms and sophisticated self-consistency alignment techniques.
@@
Visual Chain-of-thought Reasoning The integration of Chain-of-thought (CoT) [56] and its variants [69, 63, 62] has significantly enhanced LLM reasoning abilities through structured step-by-step processes. For the multimodal reasoning challenge, researchers have developed various CoT approaches [37, 70, 61, 58] that establish appropriate reasoning anchors and extend reasoning pathways. Shikra [6] improves referential expression by incorporating specific visual regions into the reasoning process, while VoCoT [27] extends visually-grounded reasoning chains. VisualCoT [46] provides a comprehensive benchmark validating CoT effectiveness in multi-hop image reasoning tasks. EmbodiedCoT [67] pioneered CoT application in embodied AI by enhancing intermediate reasoning through fine-tuning OpenVLA [21]. In contrast, FSD uniquely integrates visual-spatial reasoning using spatial relationship graphs as reasoning anchors.
@@
Visual Aids Empowered Robotic Manipulation Extracting embodiment-agnostic visual aids to enhance training efficiency has emerged as a promising paradigm in robotic manipulation. Numerous studies [1, 57, 59, 72, 65] have explored robotic policy learning based on visual traces, though most remain confined to specific tasks with cross-embodiment applicability. LLaRVA [39] advances this field by predicting visual traces to better align visual and action spaces for robot learning, compiling an impressive large visual trace dataset, yet struggles to generalize to novel downstream tasks without task-specific fine-tuning. Spatial affordance represents another effective visual aid, with several works [66, 38, 42, 48, 16, 60, 26] demonstrating its utility in robotic manipulation tasks. Our key insight is that the scarcity of embodied data fundamentally limits purely data-driven visual aid prediction approaches, impeding zero-shot generalization to unseen scenarios. Therefore, we employ a reasoning-driven paradigm to activate general spatial abilities and enhance generalization.
@@ 3.1 Definition of Visual Aids
As shown in Fig. 2, FSD utilizes three visual aids of increasing complexity, all defined within normalized image coordinates $\mathbf{x}=(p,q)\in[0,1000]^{2}\subset\mathbb{R}^{2}$: Spatial affordance boxes $\mathcal{B}=[x_{1},y_{1},x_{2},y_{2}]$ define target regions for object placement. For instructions like "Put the sushi into the silver pot," the model must infer coordinates for unmarked free space—beyond standard detector capabilities. Spatial affordance points $\mathcal{P}=\{(x_{i},y_{i})\mid i=1,2,...,n\}$ provide more precise and flexible placement with reduced redundancy. Object-centric visual traces $\boldsymbol{\tau}=\{\mathbf{x}_{t}\mid t=1,2,...,T\}$ represent ordered coordinate sequences that describe manipulation trajectories, where $T$ denotes the sequence length. These traces enable complex instruction execution, cross-embodiment transfer, and collision avoidance. We implement these representations in 2D rather than 3D due to limited high-quality 3D data availability [68]. Our object-centric rather than agent-centric visual aids approach effectively circumvents the limitations of heterogeneous embodied data while leveraging general visual datasets without robots, thus enabling robust generalization to novel scenarios and tasks.
@@ 3.2 Spatial Relationship-Focused Visual Chain-of-thought
To enable VLMs to generate spatial visual aids, a direct approach is supervised fine-tuning [39, 66, 23, 67] using these aids as a new action space or employing generative models [47, 59]. However, the heterogeneity and scarcity of embodied datasets limit this method. Directly aligning RGB images with coordinate points is challenging and prone to overfitting, hindering generalization. How can we stimulate VLMs’ spatial reasoning abilities to guide the generation rather than merely relying on extensive demonstrations? Inspired by human cognition (Fig. 3 (Top)), when executing tasks like "putting broccoli into a pot," humans first locate relevant objects, then plan movement paths based on relative positions while considering feasibility and obstacles. During this process, humans construct reasoning chains, repeatedly referencing object positions and establishing spatial relationships.
Based on these considerations, we introduce Spatial Relationship-Focused Visual Chain-of-thought (SrCoT). This approach guides VLMs to generate visual aids through structured reasoning based on spatial relationship graphs. SrCoT consists of two essential phases: 1 Description: We generate object-centric region captions establishing a task-relevant spatial relationship graph where nodes represent objects with their coordinates and edges denote relative relationships (above, below, left, right, etc.). 2 Reasoning: Using the spatial relationship graph as anchor points, we determine start and end coordinates through object references and free space reasoning, then iteratively derive intermediate points with explicit logical connections between steps. Thus, we prescribe a templated reasoning path for VLMs, enabling FSD to perform analogical reasoning in the spatial domain. While VLMs struggle to directly map future actions to image coordinates, our method leverages known object relationships as reference points for multi-hop analysis, simplifying the reasoning process. Fig. 3 (Bottom) demonstrates a complete reasoning sequence. This step-by-step SrCoT approach, though powerful, fundamentally depends on precise spatial understanding capabilities.
@@ 3.3 Weak-to-Strong Capability Dataset Construction Pipeline
The SrCoT mechanism demands enhanced capabilities from VLMs, including precise reference grounding, spatial understanding, and complex instruction-following capabilities (directly predicting future point trajectories) where mainstream VLMs [34, 30, 44, 8] show limitations. Consequently, we designed a weak-to-strong data construction pipeline to progressively develop these abilities.
FSD encompasses five hierarchical capability levels: 1 Region Grounding enables robots to focus on key objects in scenes. Although grounding capabilities have been broadly integrated into current VLMs [6, 64], understanding various small objects and complex scenes for embodied tasks is still limited; 2 Spatial Relationship understanding establishes prerequisite knowledge for spatial reasoning, forming the anchor points for SrCoT; 3 Spatial Reasoning builds upon these foundations to perform multi-hop analysis of object positions and relationships; and finally, 4 Spatial Affordance Generation and 5 Visual Trace Generation create actionable spatial guidance. Notably, SrCoT functions as a general visual-spatial reasoning mechanism applicable beyond visual traces to diverse spatial reasoning tasks. Through hierarchical spatial capability training, we enhance VLMs’ general spatial reasoning abilities, extending well beyond just embodied domains.
@@
Automatic Dataset Construction: We leveraged extensive robot data from BridgeDataV2 [53], RT-X [40], and Droid [20] to construct FSD’s training dataset. Inspired by SpatialVLM [5] and SpatialRGPT [7], our automated data construction pipeline created 300K Supervised Fine-Tuning (SFT) data across numerous formats. After filtering demonstrations with unclear instructions, we used GPT4o [15] to nominate task-relevant objects while excluding out-of-range or overly complex items. We then built around objects using GroundedSam [45] for bounding boxes and segmentation masks (Level 1 dataset). For spatial relationship labels, we reconstructed 3D semantic scene graphs using Metric3Dv2 [13] for depth estimation, along with WildCamera [73] and PerspectiveFields [17] for camera parameters. This enabled 2D-3D mapping and spatial relationship graph construction. (Level 2 dataset). We only generate relative depth sorting data to infer positional relationships, so the accuracy requirement is not high. To improve data quality, in particular, we selected objects with a relative depth gap of 20% for subsequent generation. Afterward, we randomly sample spatial relationship graphs to construct spatial reasoning QA. We provide image captions, object coordinates, and relationships as a context for GPT4o to create complex QAs (Level 3 Dataset).
A core aspect of the FSD dataset is the visual aids generation. We employ a simple method with successful human demonstrations from embodied datasets and infer the process from the results. Spatial affordance represents the designated completion area for manipulation tasks. To create spatial affordance labels (Level 4 Dataset), we extract the manipulated object’s final position from the terminal frame, combine it with reference object positioning, calculate the precise affordance region, and re-render this information onto the initial frame. For visual trace generation (Level 5 Dataset), we employ a two-stage approach: first applying self-supervised keypoint extraction [14] to identify grasp points on manipulated objects, then utilizing Cotracker [19] to capture temporal dynamics from human demonstrations, subsequently projecting these trajectories onto the initial frame. Throughout this process, we employed strict rule-based filters and continually validated our approach against manually annotated test sets, iteratively refining our filtering criteria based on empirical feedback to ensure the resulting dataset met our quality requirements. The dataset presentation, data filtering process, and prompts used to generate the data are provided in Appendix A.

### v2

@@ Abstract
Achieving generalization in robotic manipulation remains a critical challenge, particularly for unseen scenarios and novel tasks. Current Vision-Language-Action (VLA) models, while building on top of general Vision-Language Models (VLMs), still fall short of achieving robust zero-shot performance due to the scarcity and heterogeneity prevalent in embodied datasets. To address these limitations, we propose FSD (From Seeing to Doing), a novel vision-language model that generates intermediate representations through spatial relationship reasoning, providing fine-grained guidance for robotic manipulation. Our approach combines a hierarchical data construction pipeline for training with a self-consistency mechanism that aligns spatial coordinates with visual signals. Through extensive experiments, we comprehensively validated FSD’s capabilities in both “seeing” and “doing”, achieving outstanding performance across 8 benchmarks for general spatial reasoning and embodied reference abilities, as well as on our proposed more challenging benchmark VABench. We also verified zero-shot capabilities in robot manipulation, demonstrating significant performance improvements over baseline methods in both SimplerEnv and real robot settings. Experimental results show that FSD achieves 40.6% success rate in SimplerEnv and 72% success rate across 8 real-world tasks, outperforming the strongest baseline by 30%. More visualizations and datasets are available on website.
Keywords: General Robotic Manipulation, VLM Reasoning, Spatial Chain-of-Thought, Spatial VLMs
= Projects: https://embodied-fsd.github.io/
[Uncaptioned image] Code Repository: https://github.com/pickxiguapi/Embodied-FSD
= Datasets: https://huggingface.co/IffYuan
= Contact: yuanyf@tju.edu.cn
@@ 1 Introduction
A driving force behind robotics research is the pursuit of generalization: creating agents capable of versatile action across diverse robotic platforms, extending beyond familiar tasks, objects, and environments while adapting to dynamic visual inputs. Current approaches Kim et al. (2024), Brohan et al. (2023) leverage pre-trained Vision-Language Models (VLMs) and transform them into Vision-Language-Action Models (VLAs) using large-scale embodied datasets. This enables systems to interpret natural language instructions and generate robotic manipulation actions Black et al. (2024). The intention is to capitalize on the generalization capabilities of VLMs pre-trained on internet-scale data, with the hope that resulting VLAs will adapt to novel scenarios involving unseen objects and tasks. However, empirical evidence Zheng et al. (2024), Zawalski et al. (2024), Liu et al. (2024c) demonstrates that directly applying the generalization power of VLMs falls short of achieving strong zero-shot performance on completely novel tasks.
@@
We attribute the limited generalization in VLA-based systems to scarcity and heterogeneity of datasets. Robotics data remains limited compared to language and vision datasets, preventing similar scaling laws Kaplan et al. (2020), Lin et al. (2024a). Despite growth in embodied datasets O’Neill et al. (2023), their insufficient coverage and diversity prevent robust zero-shot generalization. Additionally, robotic embodiment heterogeneity Wang et al. (2024) causes significant variation in VLA outputs, making end-to-end supervised learning from vision and language to diverse action outputs a potentially unrealistic path toward generalization.
@@
We present a novel pipeline addressing generalization challenges in robotic manipulation. Our approach leverages VLMs’ visual understanding capabilities, augmented with step-by-step reasoning to extract unified mid-level structural representations independent of robot embodiment. This representation is key to generalizing learning across diverse physical interactions and dynamic behaviors. Specifically, this mid-level representation includes spatial affordance boxes/points and visual traces, each represented as marked coordinates within visual images. These visual aids provide expressive yet compact spatial information that enables effective reasoning and decisions, overcoming both scarcity and heterogeneity limitations. We introduce FSD (From Seeing to Doing), a model that generates visual intermediate representations through spatial reasoning (Fig. 1). FSD comprises three key components: (1) Spatial Relationship-Focused Visual Chain-of-Thought (SrCoT), which conducts multi-step reasoning anchored by object coordinates and spatial relationships, treating visual aid generation as a reasoning process; (2) A hierarchical data construction pipeline combining large-scale embodied datasets with common sense data, establishing a weak-to-strong capability enhancement training process; and (3) A self-consistency mechanism that aligns understanding and generation by binding spatial coordinates with specific visual signals. To evaluate the accuracy and generalization ability of visual aids generated in complex scenes, we also manually annotated 300 images from both real-world and simulation tasks across various scenarios, forming a Visual Aids Generation Benchmark (VABench). Through carefully crafted training and evaluation methods, FSD can achieve precise generation of visual aids in different scenarios, then the robot follows spatial affordances and visual traces through simple planning methods to complete action execution.
@@
FSD generalizes effectively to new instructions, and scenes through enhanced reasoning abilities. Our contributions include: (1) A novel paradigm bridging VLM reasoning with embodied decisions via visual aids; (2) The SrCoT method enabling multi-step reasoning for visual aid generation and guiding zero-shot manipulation; (3) Our weak-to-strong spatial reasoning and visual aids datasets, along with VABench, a manually annotated challenging benchmark for visual aids generation; (4) Superior performance across 8 benchmarks in spatial reasoning, free space reference, and visual aids generation, with zero-shot deployment achieving 40.6% success in SimplerEnv and 72% in 8 real-world tasks, outperforming RoboPoint baseline by 30%.
@@ 2 Related Work
Spatial Understanding and Reasoning with VLMs Spatial understanding and reasoning Liu et al. (2023b), Song et al. (2024), Du et al. (2024), Liao et al. (2024), Ray et al. (2024) require VLMs to infer spatial information beyond 2D RGB images, a capability crucial for embodied AI applications such as navigation Song et al. (2024), Hong et al. (2023), Li et al. (2024b) and manipulation Yuan et al. (2024b). Recent advances include SpatialVLM Chen et al. (2024), which explicitly incorporates spatial primitives and coordinate systems to enhance geometric reasoning capabilities. Similarly, SpatialRGPT Cheng et al. (2024) and SpatialBot Cai et al. (2024) improve spatial capabilities through more precise spatial relationship modeling. FSD specifically targets embodied manipulation scenarios by enhancing spatial reasoning capabilities through novel SrCoT mechanisms and sophisticated self-consistency alignment techniques.
@@
Visual Chain-of-thought Reasoning The integration of Chain-of-thought (CoT) Wei et al. (2022) and its variants Zhang et al. (2022), Yao et al. (2023b, 2024) has significantly enhanced LLM reasoning abilities through structured step-by-step processes. For the multimodal reasoning challenge, researchers have developed various CoT approaches Mitra et al. (2024), Zheng et al. (2023a), Yao et al. (2023a), Wu et al. (2025) that establish appropriate reasoning anchors and extend reasoning pathways. Shikra Chen et al. (2023) improves referential expression by incorporating specific visual regions into the reasoning process, while VoCoT Li et al. (2024f) extends visually-grounded reasoning chains. VisualCoT Shao et al. (2024) provides a comprehensive benchmark validating CoT effectiveness in multi-hop image reasoning tasks. EmbodiedCoT Zawalski et al. (2024) pioneered CoT application in embodied AI by enhancing intermediate reasoning through fine-tuning OpenVLA Kim et al. (2024). In contrast, FSD uniquely integrates visual-spatial reasoning using spatial relationship graphs as reasoning anchors.
@@
Visual Aids Empowered Robotic Manipulation Extracting embodiment-agnostic visual aids to enhance training efficiency has emerged as a promising paradigm in robotic manipulation. Numerous studies Bharadhwaj et al. (2024), Wen et al. (2023), Xu et al. (2024), Zheng et al. (2024), Yuan et al. (2024a) have explored robotic policy learning based on visual traces, though most remain confined to specific tasks with cross-embodiment applicability. LLaRVA Niu et al. (2024) advances this field by predicting visual traces to better align visual and action spaces for robot learning, compiling an impressive large visual trace dataset, yet struggles to generalize to novel downstream tasks without task-specific fine-tuning. Spatial affordance represents another effective visual aid, with several works Yuan et al. (2024b), Mo et al. (2021), Qin et al. (2020), Song et al. (2024), Ji et al. (2025), Yang et al. (2025), Li et al. (2025) demonstrating its utility in robotic manipulation tasks. Our key insight is that the scarcity of embodied data fundamentally limits purely data-driven visual aid prediction approaches, impeding zero-shot generalization to unseen scenarios. Therefore, we employ a reasoning-driven paradigm to activate general spatial abilities and enhance generalization.
@@ 3.1 Definition of Visual Aids
As shown in Fig. 2, FSD utilizes three visual aids of increasing complexity, all defined within normalized image coordinates $\mathbf{x}=(p,q)\in[0,1000]^{2}\subset\mathbb{R}^{2}$: Spatial affordance boxes $\mathcal{B}=[x_{1},y_{1},x_{2},y_{2}]$ define target regions for object placement. For instructions like "Put the sushi into the silver pot," the model must infer coordinates for unmarked free space—beyond standard detector capabilities. Spatial affordance points $\mathcal{P}=\{(x_{i},y_{i})\mid i=1,2,...,n\}$ provide more precise and flexible placement with reduced redundancy. Object-centric visual traces $\boldsymbol{\tau}=\{\mathbf{x}_{t}\mid t=1,2,...,T\}$ represent ordered coordinate sequences that describe manipulation trajectories, where $T$ denotes the sequence length. These traces enable complex instruction execution, cross-embodiment transfer, and collision avoidance. We implement these representations in 2D rather than 3D due to limited high-quality 3D data availability Zhang et al. (2024). Our object-centric rather than agent-centric visual aids approach effectively circumvents the limitations of heterogeneous embodied data while leveraging general visual datasets without robots, thus enabling robust generalization to novel scenarios and tasks.
@@ 3.2 Spatial Relationship-Focused Visual Chain-of-thought
To enable VLMs to generate spatial visual aids, a direct approach is supervised fine-tuning Niu et al. (2024), Yuan et al. (2024b), Li et al. (2024c), Zawalski et al. (2024) using these aids as a new action space or employing generative models Shridhar et al. (2024), Xu et al. (2024). However, the heterogeneity and scarcity of embodied datasets limit this method. Directly aligning RGB images with coordinate points is challenging and prone to overfitting, hindering generalization. How can we stimulate VLMs’ spatial reasoning abilities to guide the generation rather than merely relying on extensive demonstrations? Inspired by human cognition (Fig. 3 (Top)), when executing tasks like "putting broccoli into a pot," humans first locate relevant objects, then plan movement paths based on relative positions while considering feasibility and obstacles. During this process, humans construct reasoning chains, repeatedly referencing object positions and establishing spatial relationships.
Based on these considerations, we introduce Spatial Relationship-Focused Visual Chain-of-thought (SrCoT). This approach guides VLMs to generate visual aids through structured reasoning based on spatial relationship graphs. SrCoT consists of two essential phases: 1 Description: We generate object-centric region captions establishing a task-relevant spatial relationship graph where nodes represent objects with their coordinates and edges denote relative relationships (above, below, left, right, etc.). 2 Reasoning: Using the spatial relationship graph as anchor points, we determine start and end coordinates through object references and free space reasoning, then iteratively derive intermediate points with explicit logical connections between steps. Thus, we prescribe a templated reasoning path for VLMs, enabling FSD to perform analogical reasoning in the spatial domain. While VLMs struggle to directly map future actions to image coordinates, our method leverages known object relationships as reference points for multi-hop analysis, simplifying the reasoning process. Fig. 3 (Bottom) demonstrates a complete reasoning sequence. This step-by-step SrCoT approach, though powerful, fundamentally depends on precise spatial understanding capabilities. To improve the stability and reliability of reasoning paths and reduce model hallucinations, SrCoT requires the model to generate coordinates in a specified format and bind them to objects while performing object-centric reasoning Wang et al. (2025), Li et al. (2024f). We use <ref> to mark objects, and <point> and <box> to mark points and boxes, respectively, ensuring that each object is strictly bound to its corresponding coordinates. This explicit visual-spatial coordinate alignment enhances the FSD ’s understanding of the spatial positions and relationships of objects. All coordinates are treated as text and are discretized as integers normalized between 0 and 999.
@@ 3.3 Weak-to-Strong Capability Dataset Construction Pipeline
The SrCoT mechanism demands enhanced capabilities from VLMs, including precise reference grounding, spatial understanding, and complex instruction-following capabilities (directly predicting future point trajectories) where mainstream VLMs Liu et al. (2024b), Lin et al. (2024b), Ray et al. (2024), Du et al. (2024) show limitations. Consequently, we designed a weak-to-strong data construction pipeline to progressively develop these abilities.
FSD encompasses five hierarchical capability levels: 1 Region Grounding enables robots to focus on key objects in scenes. Although grounding capabilities have been broadly integrated into current VLMs Chen et al. (2023), You et al. (2023), understanding various small objects and complex scenes for embodied tasks is still limited; 2 Spatial Relationship understanding establishes prerequisite knowledge for spatial reasoning, forming the anchor points for SrCoT; 3 Spatial Reasoning builds upon these foundations to perform multi-hop analysis of object positions and relationships; and finally, 4 Spatial Affordance Generation and 5 Visual Trace Generation create actionable spatial guidance. Notably, SrCoT functions as a general visual-spatial reasoning mechanism applicable beyond visual traces to diverse spatial reasoning tasks. Through hierarchical spatial capability training, we enhance VLMs’ general spatial reasoning abilities, extending well beyond just embodied domains.
@@
Automatic Dataset Construction: We leveraged extensive robot data from BridgeDataV2 Walke et al. (2023), RT-X O’Neill et al. (2023), and Droid Khazatsky et al. (2024) to construct FSD’s training dataset. Inspired by SpatialVLM Chen et al. (2024) and SpatialRGPT Cheng et al. (2024), our automated data construction pipeline created 300K Supervised Fine-Tuning (SFT) data across numerous formats. After filtering demonstrations with unclear instructions, we used GPT4o Hurst et al. (2024) to nominate task-relevant objects while excluding out-of-range or overly complex items. We then built around objects using GroundedSam Ren et al. (2024) for bounding boxes and segmentation masks (Level 1 dataset). For spatial relationship labels, we reconstructed 3D semantic scene graphs using Metric3Dv2 Hu et al. (2024) for depth estimation, along with WildCamera Zhu et al. (2024) and PerspectiveFields Jin et al. (2023) for camera parameters. This enabled 2D-3D mapping and spatial relationship graph construction. (Level 2 dataset). We only generate relative depth sorting data to infer positional relationships, so the accuracy requirement is not high. To improve data quality, in particular, we selected objects with a relative depth gap of 20% for subsequent generation. Afterward, we randomly sample spatial relationship graphs to construct spatial reasoning QA. We provide image captions, object coordinates, and relationships as a context for GPT4o to create complex QAs (Level 3 Dataset).
A core aspect of the FSD dataset is the visual aids generation. We employ a simple method with successful human demonstrations from embodied datasets and infer the process from the results. Spatial affordance represents the designated completion area for manipulation tasks. To create spatial affordance labels (Level 4 Dataset), we extract the manipulated object’s final position from the terminal frame, combine it with reference object positioning, calculate the precise affordance region, and re-render this information onto the initial frame. For visual trace generation (Level 5 Dataset), we employ a two-stage approach: first applying self-supervised keypoint extraction Huang et al. (2024) to identify grasp points on manipulated objects, then utilizing Cotracker Karaev et al. (2024) to capture temporal dynamics from human demonstrations, subsequently projecting these trajectories onto the initial frame. Throughout this process, we employed strict rule-based filters and continually validated our approach against manually annotated test sets, iteratively refining our filtering criteria based on empirical feedback to ensure the resulting dataset met our quality requirements. The dataset presentation, data filtering process, and prompts used to generate the data are provided in Appendix A.